# Privacy-Preserving Dataset Combination

### v2

@@ Abstract
Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare. This dynamic especially disadvantages smaller organizations that lack resources to purchase data or negotiate favorable sharing agreements. We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information. Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving $>90\%$ correlation with plaintext evaluations. In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy. Our framework provides a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. These results demonstrate the potential for privacy-preserving data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data resources. Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
@@ 1 Introduction
Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size [1, 2]. These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups [13, 14].
@@
However, access to data varies significantly across entities and domains. While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access. Domain-specific data is also becoming increasingly valuable for fine-tuning these general models [15, 16, 17], creating a competitive advantage for the data owners. As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets [18, 19, 20]. This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
Scaling datasets in regulated domains is particularly challenging due to legal constraints and unpredictable outcomes from altering the training data composition. In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy. Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it [21, 22, 23]. Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources [24, 25, 13]. This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets [13, 26]. However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it. This reluctance to share creates a bottleneck in improving model performance. In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes [27, 28, 29, 30].
To address these challenges, we propose SecureKL, a privacy-preserving approach for guiding dataset combinations without requiring direct data or model sharing. Our work makes three primary technical contributions:
  1. 1.
    We introduce a framework for practical secure data combination that enables organizations to evaluate potential partnerships while maintaining data privacy. Our framework categorizes existing approaches by their privacy leakage risks and provides a systematic way to assess the trade-offs between data utility and privacy preservation.
  2. 2.
    We extend KL-XY score [13] with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships. Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving $>90\%$ correlation with plaintext computations.
  3. 3.
    Through extensive evaluation in low-data, high-stakes settings, we demonstrate that our method successfully identifies beneficial data partnerships for intensive care unit (ICU) mortality prediction, improving classifier performance for the source hospital.
  4. 4.
    In experimental scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
We argue that our method presents an appealing trade-off of privacy and utility by preserving privacy for both parties while using all the underlying data.
@@ 2 Problem Setup
$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline Performance)
@@
$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined Performance)
We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$. To add a single additional data source by setting $T=\{i\}$, the improvement is $\delta_{i}=\delta_{(o,i)}=\mathrm{AUC}_{i}-\mathrm{AUC}_{o}$. This leads to our central question: Without seeing target data, how does a hospital ascertain potential data sources to combine with?
@@
$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$ (Ideal Dataset Combination)
@@
Computing every subset $T\subset{[1...N]\choose n}$’s associated $\delta_{T}$ is exponential in $n$. To make this problem tractable, we make two key assumptions. First, we apply strategies greedily, selecting top-ranked target datasets. With the ultimate objective of improving the source hospital’s prediction task, we fix $H_{o}$; to compare the trade-offs between strategies in Section 3, we apply each $\pi$ greedily to select top-$n$ institution(s) for $H_{o}$ without replacement. Second, in in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
@@
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work [13]. KL divergence [31], also called information gain [32], describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (Kullback–Leibler Divergence)
Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  [13] proposes two groups of scores to make this divergence approximation tractable from small samples. Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership. Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY Score)
Details are described in Section 3.
Privacy Model for $\pi_{p}$.
@@
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon methods in  [13] .
MPC Preliminary
To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) [33, 34] protocols are leveraged. Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties. This is implemented with the research framework CrypTen [35], specialized for MPC and machine learning. Our algorithmic and engineering details are in Sections  3 and  4, respectively. For related secure techniques, see Section 5.4.
Additional Assumptions
@@
Generally, we consider high stakes domains where disparate data may have additive benefits to the existing data. In order to make privacy boundaries tractable, we make the following additional assumptions:
@@
    Legal risks of sharing any data are omnipresent in high stakes domains. The risks with sharing sensitive data in $\pi_{d}$ and $\pi_{s}$ are not made explicit, but assumed to be ”medium” and ”medium-to-high” respectively. This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
@@ 3 Methods
Figure 1: Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$. Each side encrypts their data, and a model is trained on their joint data. Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
We summarize our data acquisition strategies differentiated by leakage risks, which correlate with potential costs:
Category 1, medium-to-high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered [36]. Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance uses  [13] ’s KL-XY Score.
Category 2, medium leakage, sharing summary statistics. $\pi_{d}(n)$ uses demographic metadata to guide data selection. This is implemented through ratio distance between source and target distributions, which may be considered aggregates therefore potentially not sensitive, such as when the underlying aggregation function $\phi$ is differentially private.
Category 3, zero leakage, sharing no additional information besides what is assumed public. There are two methods: a. Blind selection baseline: $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. Prior works suggests that when $n=1$, randomly selecting a source in hospital ICU may be risky and inefficient. b. Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private KL-XY.
3.1 Trivial Baseline: Blind Selection
Blind selection refers to the process when no information is provided. $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  [13] suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
3.2 Sharing Summary Statistics
3.3 KL-based Methods, in Plaintext
$\pi_{s}(n,k)$ assumes each of the candidate hospitals will share a set of raw data. In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30.300,3000\}$ (Section 4.3). Though leakage can be controlled through $k$, the data is inherently sensitive.
This is implemented with KL-based measures similar to [13] . To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:

### v3

@@ Abstract
Privacy concerns and competitive interests impede data access for machine learning, due to the inability to privately assess external data’s utility. This dynamic disadvantages smaller organizations that lack resources to aggressively pursue data-sharing agreements. In data-limited scenarios, not all external data is beneficial, and collaborations suffer especially in heavily regulated domains: metrics that aim to assess external data given a source e.g., approximating their KL-divergence, require accessing samples from both entities pre-collaboration, hence violating privacy (Compton et al., 2023; Shen et al., 2024). This conundrum disempowers legitimate data-sharing, leading to a false “privacy-utility trade-off”. To resolve privacy and uncertainty tensions simultaneously, we introduce SecureKL, the first secure protocol for dataset-to-dataset evaluations with zero privacy leakage, designed to be applied preceding data sharing. SecureKL evaluates a source dataset against candidates, performing dataset divergence metrics internally with private computations, all without assuming downstream models. On real-world data, SecureKL achieves high consistency ($90\%$ correlation with non-private counterparts) and successfully identifies beneficial data collaborations in highly-heterogeneous domains (ICU mortality prediction across hospitals and income prediction across states). Our results highlight that secure computation maximizes data utilization, outperforming privacy-agnostic utility assessments that leak information.
@@ 1 Introduction
Figure 1: Privacy can dis-incentivize data collaborations. Without seeing external data, an organization has two strategies: i. blind default $\pi_{0}$: randomly selecting partnerships causes hesitation and hinders partnerships. ii. private evaluation $\pi_{p}$: securely assessing datasets before commitment.
Since Hestness et al. (2017), empirical works have predetermined data as a key driver to performance gains, via the so-called “scaling laws” (Kaplan et al., 2020; Brown et al., 2020). Yet, accessing and combining datasets is persistently challenging. As datasets have evolved from small to larger (Sun et al., 2017), more diverse (Zhang et al., 2020), and more compute-optimal (Hoffmann et al., 2022), the field of machine learning continues to seek more data  (Mahajan et al., 2018; Brown et al., 2020; Park et al., 2019; Reed et al., 2022; Sheller et al., 2020) and better ways to combine it  (Shen et al., 2023; Nguyen et al., 2023; Petty et al., 2024; Li et al., 2024; McKinney et al., 2020).
@@
Strategically combining data from different sources promises enhanced models, but disempowers smaller organizations. While diverse, high quality data often improves performance, robustness, and fairness (Miller et al., 2021), access to such data significantly varies across entities and domains (Tenopir et al., 2011; L’Heureux et al., 2017). As domain-specific data becomes increasingly valuable (Alsentzer et al., 2019; Gururangan et al., 2020; Lee et al., 2020), data-owning entities are more reluctant to share it for free, opting instead to sell it in emerging data markets (Acemoglu et al., 2022; Huang et al., 2021; Liang et al., 2018). This dynamic disproportionately handicaps smaller organizations who lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
Organizations may hesitate to commit to a potential partnership when unsure about the benefits. As Figure 1 illustrates, this “commitment issue” is not solely a privacy issue; it’s the inability to privately assess an external dataset’s utility before partnerships. This evaluation is, however, nontrivial. For example, adding in-domain data does not necessarily result in more performant models. Due to the precarious nature of domain shifts, machine learning models are sensitive to additional training data, often unpredictably when the source dataset is small (e.g., a single hospital’s data) – a phenomenon known as “the dataset combination problem” (Koh et al., 2021; Wang et al., 2022; Taori et al., 2020; Bradley et al., 2021; Meng, 2018; Shen et al., 2024).
Figure 2: The Dataset Combination Problem. Real-world data collaborations are inherently uncertain. AUC change for a source entity, after incorporating external data across hospitals and states. Each entry representas training on a pairwise data combination for a test entity (x-axis) Each column represents training on a data combination for a fixed test dataset, formed by pairing the fixed entity with an external dataset in the y-axis. Left: In eICU (Pollard et al., 2018), 10 out of 12 hospitals may see their mortality prediction model degrade for some potential hospital partners. Right: In Folktables (Ding et al., 2021), combining with random state leads to worse income prediction in 10 out of 15 states. (red is bad; exact values are reported in Appendix 15)
Ideally, all target data should be considered to reduce uncertainty in costly data collaborations. Yet, datasets owned by separate entities cannot be directly and fully accessed, significantly limiting the practicality of non-private dataset measures (Shen et al., 2024; Ilyas et al., 2022).
Our work directly addresses this crux by recognizing both privacy and competitive incentives. First, before committing to acquiring unseen data, we enable organizations to privately gauge the relative utility of candidate datasets. Second, we provide strong privacy guarantees required of entities operating under stringent regulations e.g. healthcare providers to navigate data acquisition.
Developers are often uncertain about the most effective model before more data becomes available. This renders a secure data appraisal stage by Xu et al. (2022), which requires model parameters, not applicable. In this opportunistic setting, we ask:
Can we ascertain the differential utility of prospective datasets, without knowing the final model?
We introduce SecureKL, the first divergence-based dataset measure with zero privacy leakage. Our key insight is that private divergence computations (via secure multiparty computation (MPC)) are more data-efficient than sharing samples, thus achieving high performance. By preserving performance, SecureKL presents a compelling guarantee: for both parties, privacy is fully protected while data utilization is maximized.
Contributions
A novel secure dataset-to-dataset evaluation protocol SecureKL (SKL) that reduces uncertainty in data utility under limited data and budget, producing privacy-preserving measures while using the maximum available samples. SKL achieves a $90\%$ correlation with privacy-violating counterparts across two real-world heterogeneous domains. Empirically, on ICU mortality prediction, SKL reliably chooses beneficial hospital(s) to partner with, outperforming data-leaking alternatives, including using demographic summaries or sharing data subsets.
Impact
We provide a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. Our major advantage lies in reliability, especially when small organizations cannot afford to invest in detrimental partnerships. These results demonstrate the potential for wider data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data. Our code is available, and can be readily deployed to demonstrate potential data value preceding collaborations [^1] https://github.com/kere-nel/secure-data-eval.
2 Background
2.1 Dataset Combination Problem
In high-stakes domains, incorporating additional datasets may degrade the model. In healthcare scenarios, both Compton et al. (2023) and Shen et al. (2024) showed that blindly acquiring new datasets can degrade model performance. We extend these findings by replicating the effect under our own experimental setup. As shown in Figure 2, combining data from one hospital or state with an external source often leads to inconsistent, and sometimes negative, changes in model performance. This non-monotonic behavior highlights the need to evaluate data partners before embarking on a full-fledged collaboration.
2.2 Trading Off Sample Utility For “Privacy”
Figure 3: Non-private evaluation strategies. iii. sub sampling $\pi_{s}$: a subset of the target’s data is shared. iv. demographic summaries $\pi_{d}$: the target entity discloses distributions by protected attributes, i.e. age, gender, or race.
Before an agreement, informed negotiations become impossible when entities do not expose data. In practice, data owners frequently resort to tiny samples (Centers for Medicare & Medicaid Services, 2025) or summary statistics (e.g., race, gender, age) for making decisions (Figure 3). Yet, in data-limited settings, model performance is highly sensitive to new input. These heuristics are fickle, as sparse traits or limited samples often fail to capture the entire dataset’s nuance and complexity, especially in heterogeneous domains, creating a perceived privacy-utility trade-off. We will show that secure computation, on the other hand, can avoid this tradeoff by measuring over entire datasets while maintaining privacy.
@@ 3 Problem Setup
$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline AUC)
@@
$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined AUC)
We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$. To add a single additional data source by setting $T=\{i\}$, the improvement is $\delta_{i}=\delta_{(o,i)}=\mathrm{AUC}_{i}-\mathrm{AUC}_{o}$. This leads to our central question:
Without seeing target data, how does a hospital ascertain potential data sources to combine with?
@@
$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset\binom{[1...N]}{n}}\mathrm{AUC}_{T}$ (Ideal Combination)
@@
Computing every subset $T\subset\binom{[1...N]}{n}$’s associated $\delta_{T}$ is exponential in $n$. To make this problem tractable, we make two key assumptions. First, we apply strategies greedily, selecting top-ranked target datasets. With the ultimate objective of improving the source hospital’s prediction task, we fix $H_{o}$; to compare the trade-offs between strategies in Section 4, we apply each $\pi$ greedily to select top-$n$ institution(s) for $H_{o}$ without replacement. Second, in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
@@
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (Shen et al., 2024). KL divergence (Kullback and Leibler, 1951), also called information gain (Quinlan, 1986), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (KL-Divergence)
Because computing KL-divergence on hospital datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial due to the high dimensionality of the data, Shen et al. (2024) proposes two groups of scores to make this divergence approximation tractable from small samples. We adopt one such score, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, which trains a classifier to distinguish between source and target samples and uses the classifier’s output probabilities to estimate dataset divergence. We describe the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score in detail in Section 4.2.
Privacy Model
@@
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon Shen et al. (2024) and  Ilyas et al. (2022).
Dataset Divergence and Downstream Utility
Building on the finding of Miller et al. (2021) where a model’s in-distribution performance is related to its out-of-distribution performance (across all model choices), we confer that dataset divergence does predict downstream model’s performance after combining the data (Table 3), using metrics introduced in Shen et al. (2024) (Section 4.2). Intuitively, in-distribution quality is paramount in low-data settings, where dataset divergence can capture greater complexity and nuance than accessing a few traits.Yet, privacy is unresolved: divergence measures entail accessing both entities’ data (Shen et al., 2024; Ilyas et al., 2022), posing significant risks for heavily-regulated entities who are liable for any data exposure (Centers for Medicare & Medicaid Services, 1996; Gervasi et al., 2022).
Secure Multiparty Computation (MPC)
To cryptographically secure our divergence computation, we use Secure Multiparty Computation (MPC)  (Yao, 1982; Shamir, 1979). MPC lets two or more parties to compute a function over private inputs, revealing only the final output (Goldwasser and Micali, 2019).
Engineering machine learning workflows in private faces differs drastically from non-private machine learning engineering, due to 1. precision configurations and normalizations (to avoid blow-ups and underflows), 2. error control and performance engineering for machine learning in bit-limited spaces, and 3. debugging, hyperparameter tuning. Nevertheless, they are provably secure.
Despite their non-trivial engineering, MPC programs enjoy strong security guarantees and relative ease of deployment. Even small organizations can deploy MPC without any specialized hardware. Thus, the algorithms developed and shared in SecureKL readily enable zero-leakage dataset-to-dataset evaluations before sharing data.
Assumptions
@@
Consider high stakes domains where disparate data may have additive benefits to the existing data. To make privacy boundaries tractable, we make the following assumptions:
@@
    Legal risks of sharing any data are omnipresent in high stakes domains. The risks with sharing sensitive data in data-leaking strategies, which we coin as $\pi_{d}$ (demographic distance) and $\pi_{s}$ (small sample), are not made explicit, but assumed to be ”moderate” and ”high” respectively. This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
@@ 4 Methods
Strategy Sub-strategies Leakage
$\pi_{0}$ Blind (baseline) n/a zero
$\pi_{p}$ Private (SecureKL) n/a minimal
$\pi_{d}$ Demographic sex, gender, race moderate
$\pi_{s}$ Subset Sampling 1%, 10%, 100% high
Table 1: Concrete selection strategies $\pi$, differentiated by leakage risks. For a strategy $p$, a set of targets is chosen, i.e., $T\leftarrow\pi(\mathcal{D}_{o},n)$.
4.1 Defining Dataset Acquisition Strategies
Given the private and non-private data acquisition strategies illustrated in Figures 1 and 3, we associate privacy leakage as a primary cost of data partnerships. This section formalizes them by their leakage risks, summarized in Table 1.
A, high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (Centers for Medicare & Medicaid Services, 2025). Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance metric follows the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score introduced by Shen et al. (2024) , which we define in the following subsection.
B, moderate leakage, sharing summary statistics. $\pi_{d}(n)$ uses demographic metadata to guide data selection. This is implemented through ratio distance between source and target distributions, which may be considered aggregates therefore potentially not sensitive, such as when the underlying aggregation function $\phi$ is differentially private.
C, zero/minimal leakage, sharing no additional information besides what is assumed public, and what our method outputs, such as a score or a ranking. There are two methods: a. Blind selection baseline: $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. Prior works suggests that when $n=1$, randomly selecting a source in hospital ICU may be risky and inefficient. b. Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
4.2 KL-based Methods, Without Privacy
In subset-sampling strategy $\pi_{s}(n,k)$, each of the candidate entities will leak a set of raw data. $\pi_{s}$ is implemented with KL-based measures similar to Shen et al. (2024). Recall $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a “metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be “close” to the target distribution $P_{i}$, but not the other way around: